Find f'(x) iff(x) = ex(1+Inx)

f(x) = x3e1/x

f(x) = √[x2-1](x+1) / (x2+1)3/2

1
Expert's answer
2021-09-09T14:29:28-0400
  • Solution: "f(x)=e^x (1+ln(x))"

By the Product Rule, we have


"f'(x)=\frac{d}{dx}(e^x(1+ln(x)))= (1+ln(x))\frac{d}{dx}e^x""+e^x \frac{d}{dx}(1+ln(x))=\{the\space Sum \space Rule\}""=(1+ln(x))\frac{d}{dx}e^x+e^x (\frac{d}{dx}1+\frac{d}{dx}ln(x))""=(1+ln(x))e^x+e^x (0+\frac{1}{x})=e^x(1+ln(x)+\frac{1}{x})"

Answer: "f'(x)=e^x (1+ln(x)+\frac{1}{x})"


  • Solution: "f(x)=x^3e^{1\/x}"

By the Product Rule, we have


"f'(x)=\frac{d}{dx}(x^3e^{1/x} )=e^{1/x}\frac{d}{dx}x^3+x^3 \frac{d}{dx}e^{1/x}"

By the Chain Rule: let u=1/x


"f'(x)=e^{1/x}\frac{d}{dx}x^3+x^3 \frac{d}{du}e^u\frac{d}{dx}(1/x)""=e^{1\/x}3x^2+x^3 e^{1\/x}(-1\/x^2)=xe^{1\/x}(3x-1)"

Answer: "f'(x)=xe^{1\/x}(3x-1)"


  • Solution: "f(x)=\frac{(x+1)\sqrt{x^2 -1}}{(x^2+1)^{3/2}}"

By the Quotient Rule, we have


"f'(x)=\frac{d}{dx}\Big(\frac{(x+1)\sqrt{x^2 -1}}{(x^2+1)^{3/2}} \Big)""=\frac{(x^2+1)^{3/2}\frac{d}{dx}((x+1)\sqrt{x^2 -1})-(x+1)\sqrt{x^2 -1}\frac{d}{dx}(x^2+1)^{3/2}}{((x^2+1)^{3/2})^2} \space (1)"

By the Product Rule


"\frac{d}{dx}((x+1)\sqrt{x^2 -1})=\sqrt{x^2 -1}\frac{d}{dx}(x+1)""+(x+1)\frac{d}{dx}\sqrt{x^2 -1} \qquad (2)"

By the Sum Rule


"\frac{d}{dx}(x+1)=\frac{d}{dx}x+\frac{d}{dx}1=1+0=1 \qquad (3)"

By the Chain Rule: let "u=x^2-1"


"\frac{d}{dx}\sqrt{x^2 -1}=\frac{d}{du}\sqrt{u}\frac{d}{dx}(x^2 -1)=\{the\space Difference \space Rule\}""=\frac{d}{du}(\sqrt{u})\Big(\frac{d}{dx}x^2 -\frac{d}{dx}1\Big)=\frac{1}{2\sqrt{x^2 -1}}(2x-0)""=\frac{x}{\sqrt{x^2 -1}}\qquad(4)"

substitute (3), (4) into (2)


"\frac{d}{dx}((x+1)\sqrt{x^2 -1})=\sqrt{x^2 -1}+\frac{x(x+1)}{\sqrt{x^2 -1}}\qquad (5)"

By the Chain Rule: let "u=x^2+1"

"\frac{d}{dx}(x^2+1)^{3/2}=\frac{du^{3/2}}{du}\frac{d}{dx}(x^2+1)=\frac{du^{3/2}}{du}(\frac{d}{dx}x^2+\frac{d}{dx}1)""=\frac 3 2 \sqrt{x^2 +1}(2x+0)=3x\sqrt{x^2 +1}\qquad (6)"

substitute (5), (6) into (1)


"f'(x)=\frac{(x^2+1)^{3/2}(\sqrt{x^2 -1}+\frac{x(x+1)}{\sqrt{x^2 -1}})-3x(x+1)\sqrt{x^2 -1}\sqrt{x^2 +1}}{(x^2+1)^3}""=\frac{\sqrt{x^2 -1}+\frac{x(x+1)}{\sqrt{x^2 -1}}}{(x^2+1)^{3/2}}-\frac{3x(x+1)\sqrt{x^2 -1}}{(x^2+1)^{5/2}}""=\frac{\sqrt{x^2 -1}}{(x^2+1)^{3/2}}+\frac{x(x+1)}{\sqrt{x^2 -1}(x^2+1)^{3/2}}-\frac{3x(x+1)\sqrt{x^2 -1}}{(x^2+1)^{5/2}}"

Answer: "f'(x)=\frac{\sqrt{x^2 -1}}{(x^2+1)^{3/2}}+\frac{x(x+1)}{\sqrt{x^2 -1}(x^2+1)^{3/2}}-\frac{3x(x+1)\sqrt{x^2 -1}}{(x^2+1)^{5/2}}"